translate the following into an algebraic expression and find the number.
"the sum of five times a number and 11 is four less than the product of eight and the number."

Answer by mananth(16946)

"the sum of five times a number and 11 is four less than the product of eight and the number."
Let the number be x
five times the number = 5x
"and 11" means add 11 to 5x
= 5x+11
product of 8 and the number = 8*x=8x
4 less
means 8x-4
They are equal
5x+11=8x-4
5x-8x=-4-11
-3x=-15
x= -15/-3
x=5